If 40 ml of a 10% methanol solution was mixed with 60 ml of a 20% solution, what would the resulting solution's percent methanol be?

Answer by Boreal(15235)

Amount of methanol total is 100 ml
40(0.1)+60(0.2)=100(x), where x is the decimal amount of methanol
4+12=100x=16
Therefore, x is 16/100 or 0.16 or 16%.
The check is that 60% of the solution is 20%, so the amount is 60% of the way from 10 to 20%.